Square root 2y+1=5
The solution for y =

Answer by checkley77(12844)

sqrt(2y+1)=5
square both sides.
4y^2+4y+1=25
4y^2+4y+1-25=0
4y^2+4y-24=0
4(y^2+y-6)=0
4(y+3)(y-2)=0
y+3=0
y=-3 ans
y-2=0
y=2 ans.